OpenStudy (faiqraees):

Use proof by contradiction to prove that there doesnt exist a smallest positive rational number

OpenStudy (alivejeremy):

What?

OpenStudy (faiqraees):

@ganeshie8

OpenStudy (mww):

Hmm you wrote rational integer. This is obsolete, integers are always rational. What is the full question?

OpenStudy (faiqraees):

Thanks for notifying

OpenStudy (faiqraees):

@mathmate

OpenStudy (faiqraees):

I am aware of the basic idea that we have to assume a smallest number in form of a/b and show that a/2b is smaller than a/b thus contradicting the statement. My question is how should I provide it formally?

OpenStudy (unklerhaukus):

if x is the smallest rational number x/2 is a smaller rational number

OpenStudy (faiqraees):

One more question Verify that 5^5 =1(mod)11. Hence find the remainder obtained on dividing 5^2016 by 11

OpenStudy (unklerhaukus):

wait, is zero a rational number?

OpenStudy (faiqraees):

No

OpenStudy (unklerhaukus):

why not?

OpenStudy (faiqraees):

Yes yes it is rational sorry

OpenStudy (unklerhaukus):

does smallest mean negative numbers then?

OpenStudy (faiqraees):

No question says positive rational numbers

OpenStudy (unklerhaukus):

now it does.

OpenStudy (faiqraees):

Can you explain the second question

OpenStudy (faiqraees):

Verify that 5^5 =1(mod)11. Hence find the remainder obtained on dividing 5^2016 by 11

OpenStudy (mww):

Better to show a/b < ka/b for all positive rational k (to keep ka/b rational) to generate the set of all positive rational numbers. Define a/b to be the smallest rational positive number, the set of positive rationals is then generated by ka/b where k is a positive rational number. Then we say a/b < ka/b as a/b is the smallest rational. We can rewrite this as 1 < k. In other words k > 1. However we defined the set k to be all positive rationals and thus k > 1 is inadequate to describe the set.

OpenStudy (mathmate):

hint: \(a^k~ mod~ n \equiv (a~ mod~ n)^k~ mod~ n\)

OpenStudy (faiqraees):

@mathmate Can you provide a proof for that, since I haven't been taught that equation in college?

OpenStudy (mathmath333):

use fermat little theorem for 5^2016 by 11

OpenStudy (faiqraees):

Can it not be solved using induction, series or summation because this question was from that chapter

OpenStudy (mathmate):

If mathematical induction is your topic, then you can start by proving (a mod n)(b mod n) \(\equiv\) ab mod n (see for example http://math.stackexchange.com/questions/104721/proving-properties-of-modular-arithmetic-by-induction ) then decompose 5^2016 into 5*(5^5)^403 and proceed with induction proof to show that 5^2016 \(\equiv\) 5(5^5) \(\equiv\)...

OpenStudy (mathmate):

or better show, use induction to show that \(a^k~ mod~ n \equiv (a~ mod~ n)^k~ mod~ n\)